OpenStudy (jennychan12):

Use the Midpoint Rule, with n = 4, to approximate the area bounded by f(x) = x^3-10x^2+25x and the x-axis between x = 0 and x = 5.

OpenStudy (jennychan12):

i found that delta x = 5/4, but i kinda forgot how to use the midpoint rule.

OpenStudy (anonymous):

ok each rectangle (delta x) = (b-a)/n = (5-0)/4 = 5/4 long now you need to find the height of each rectangle at its point. So find the height of the first rectangle at (5/4)/2 = 5/8 by plugging in 5/8 into f(x) and multiply that by delta x. Do the same for the 4 rectangles. The next midpoint will be at x= 5/4 + 5/8 = 15/8

OpenStudy (anonymous):

sorry meant to say "now you need to find the height of each rectangle at its midpoint." in the second line

OpenStudy (jennychan12):

why do you divide 5/4 by 2? is it because it's midpoint?

OpenStudy (anonymous):



OpenStudy (anonymous):

Exactly, the midpoint is halfway to 5/4, so divide it by 2

OpenStudy (jennychan12):

ohhh okay, thanks :)

OpenStudy (anonymous):

For the subsequent midpoints, just add 5/8 to whatever the starting point of the rectangle is

OpenStudy (jennychan12):

ok, thanks. my math teacher didn't really "explain" this stuff :/